### v1

# FlashBias: Fast Computation of Attention with Bias

Abstract
Attention mechanism has emerged as a foundation module of modern deep learning models and has also empowered many milestones in various domains. Moreover, FlashAttention with IO-aware speedup resolves the efficiency issue of standard attention, further promoting its practicality. Beyond canonical attention, attention with bias also widely exists, such as relative position bias in vision and language models and pair representation bias in AlphaFold. In these works, prior knowledge is introduced as an additive bias term of attention weights to guide the learning process, which has been proven essential for model performance. Surprisingly, despite the common usage of attention with bias, its targeted efficiency optimization is still absent, which seriously hinders its wide applications in complex tasks. Diving into the computation of FlashAttention, we prove that its optimal efficiency is determined by the rank of the attention weight matrix. Inspired by this theoretical result, this paper presents FlashBias based on the low-rank compressed sensing theory, which can provide fast-exact computation for many widely used attention biases and a fast-accurate approximation for biases in general formalization. FlashBias can fully take advantage of the extremely optimized matrix multiplication operation in modern GPUs, achieving 1.5$\times$ speedup for AlphaFold, and over 2$\times$ speedup for attention with bias in vision and language models without loss of accuracy.
1 Introduction
In recent years, Transformer NIPS2017_3f5ee243 has enabled impressive achievements in extensive areas, including computer vision liu2021Swin ; zhao2021point , natural language processing achiam2023gpt ; touvron2023llama , scientific applications abramson2024accurate ; jumper2021highly ; wang2023scientific , etc. Especially, as the core design of Transformer, the attention mechanism with powerful relation modeling capacity has emerged as a foundation module in deep learning models, making its optimization of vital importance. As a seminal progress, FlashAttention dao2022flashattention ; dao2023flashattention speeds up the computation of standard attention by successfully reducing the read-write cost of GPU high bandwidth memory (HBM) with IO-aware techniques. Afterwards, researchers further extend FlashAttention to support diverse kinds of masks sharma2024efficiently ; wang2024flashmask , such as causal mask in decoder-only Transformers or sliding window mask.
While standard attention or attention with masks has enjoyed elaborative efficiency optimization, we notice that attention with bias is a similarly important extension in dealing with complex tasks, where useful prior knowledge is introduced as a bias term of dot-product attention weights to guide the model learning. For example, in vision and language Transformers, the relative distance among tokens is a well-established attention bias and has been proven essential to the final performance of various tasks liu2021Swin ; liu2022swin ; press2022train . Also, for scientific problems with rich domain-specific prior wang2023scientific , attention bias is an indispensable component, such as the pair representation bias in AlphaFold abramson2024accurate ; jumper2021highly ; senior2020improved . However, the targeted efficiency optimization for attention with bias is still lacking. All previous research in extending standard attention is centered on the speedup of attention with masks sharma2024efficiently ; wang2024flashmask , where the sparsity nature of masks enables possible computation reduction. However, unlike the attention mask that defines computation logic, the bias matrix describes the pairwise relation among tokens, which is inherently continuous and dense as showcased in Figure 1, making previous sparsity-based speedup techniques inapplicable. Although FlexAttention dong2024flex , benefiting from compiler techniques, a new feature in PyTorch 2.5 Paszke2019PyTorchAI , can support general formalizations of bias terms, it still depends on element-wise operations that are less optimized than matrix multiplications and fails in speeding up dynamic bias. To date, the fast computation of attention with bias remains a nascent area to explore.
Figure 1: (a-b) Comparison between attention mask and bias, where the spatial bias is from Swin Transformer liu2021Swin for computer vision, ALiBi bias is used in language modeling press2022train and pair bias is from AlphaFold abramson2024accurate . (c) FlashAttention needs to read bias tensors from HBM to the on-chip SRAM.
In FlashAttention dao2022flashattention , researchers find that instead of computation FLOPS, the read-write (IO) overload of GPU high bandwidth memory (HBM) is the actual bottleneck of speed. As bias matrices are usually dense, it is really hard to bypass the quadratic IO complexity, where the computation needs to read the whole bias term from HBM at least once, making its speedup intractable. In this paper, we notice that the IO challenge that we face here can be recast into the classical compressed sensing problem donoho2006compressed , whose basic assumption is that the “measurement” (corresponding to IO overload here) is expensive but the computation (corresponding to the fast on-chip computation) is cheap. This new perspective offers us valuable theoretical understandings. Specifically, for a dense matrix, the optimal “measurement” (IO overload) is highly related to the matrix’s rank candes2010power . This inspires us to dive into the rank of attention bias and surprisingly find that most of the widely used biases are inherently of low rank, eventually discovering one flexible way to enable fast computation of attention with bias.
Based on the above understandings, this paper presents FlashBias based on the low-rank compressed sensing theory. By formalizing commonly used attention biases into the multiplication of two factor tensors, FlashBias achieves fast and exact computation for many Transformers, covering vision, language and scientific domains. Going further, we present an approximation method for general bias formalization with a low-rank neural decomposition technique, which successfully speeds up more complicated attentions in AlphaFold 3 abramson2024accurate . Our contributions are summarized as follows:
  • Based on an in-depth study of the computation bottleneck of attention with bias, this paper theoretically proves that for a dense matrix, e.g. dot-product attention weights or various attention biases, the optimal efficiency in GPU is determined by the rank of the matrix.
  • Inspired by theoretical analyses, this paper presents FlashBias based on the low-rank compressed sensing theory, which utilizes exact and SVD decompositions for fast computation of many widely-used attention biases, and an approximation version for general biases.
  • FlashBias can accelerate a family of widely-used backbones without loss of accuracy, which brings 1.5$\times$ speedup for AlphaFold and over 2$\times$ speedup for vision and language models.
2 Preliminaries
2.1 Attention with Bias
Attention NIPS2017_3f5ee243 contains queries $\mathbf{q}\in\mathbb{R}^{N\times C}$, keys $\mathbf{k}\in\mathbb{R}^{M\times C}$ and values $\mathbf{v}\in\mathbb{R}^{M\times C}$, where $N,M$ denote the sequence length of queries and keys respectively and $C$ represents the channel of hidden representations. Conventionally, attention weights are calculated based on the dot-product between queries and keys. Beyond solely relying on the representation dot-product, useful prior knowledge is also commonly introduced into the attention mechanism as a bias term to guide learning, which is:
$\begin{split}\mathbf{o}=\operatorname{softmax}(\frac{\mathbf{q}\mathbf{k}^{\top}}{\sqrt{C}}+\mathbf{b})\mathbf{v}.\end{split}$ (1)
Here $\mathbf{o}\in\mathbb{R}^{N\times C}$ denotes results and $\mathbf{b}\in\mathbb{R}^{N\times M}$ represents the prior weight for query-key pairs, which is also referred to as attention bias jumper2021highly ; liu2021Swin . Actually, attention bias has been widely used in diverse domains and has been proven to be indispensable for model learning, especially for complex tasks. For example, in computer vision, Swin Transformer liu2021Swin adopts the relative distance among different pixels as attention biases to introduce the essential locality inductive bias into vision backbones. Similarly, relative position is also used to encode sequential information in language models glm2024chatglm ; touvron2023llama . In addition, attention bias widely exists in Transformer-based scientific deep models to incorporate domain-specific priors, such as the pair representation bias in AlphaFold abramson2024accurate ; jumper2021highly ; senior2020improved .
It is worth noting that attention mask is also commonly used in attention mechanisms to enable the calculation under a specific logic, such as the upper triangle mask (aka. causal mask) in decoder-only Transformers achiam2023gpt and sliding window mask in sparse Transformers beltagy2020longformer . Usually, the mask is also applied as an additive term to the query-key dot-product. However, attention biases that we focus on in this paper are usually dense continuous values that incorporate detailed prior knowledge into each query-key pair to reflect the degree of token relations, while masks only contain zeros and negative infinite values and are usually sparse. This difference makes their speed-up strategies distinct.
2.2 Fast Computation of Attention
As the core component in modern deep models achiam2023gpt ; jumper2021highly ; liu2021Swin ; peebles2023scalable , the attention mechanism has been widely explored. Although it has demonstrated impressive performance, one serious drawback is its computational complexity, which is quadratic w.r.t. sequence length. To tackle the efficiency bottleneck, the fast and exact computation of the attention mechanism has also been widely investigated. Based on tiling techniques, researchers formalized the softmax function in attention with a sequential calculation process milakov2018online ; rabe2021self , successfully reducing memory cost to linear complexity. Afterwards, FlashAttention dao2022flashattention and FlashAttention-2 dao2023flashattention further demystify that the IO overload of GPU HBM is the actual efficiency bottleneck and enable significant speedup by utilizing the super-fast on-chip SRAM. The FlashAttention family has served as a default acceleration option of modern Transformers and has been built-in and supported by PyTorch Paszke2019PyTorchAI .
Further, to extend the model’s capability in handling diverse types of attention, researchers have made a great effort to enable fast computation for attention with masks, such as the causal mask achiam2023gpt and sliding window mask beltagy2020longformer , etc. Technically, the design principle in speeding up attention with masks is to utilize the sparsity in masks to reduce the IO complexity. For instance, Binary Block Masking sharma2024efficiently splits the mask matrix as blocks and adopts binary values to indicate the blocks with non-zero masks, which enables the masking process to skip the IO read of all-zero mask blocks. Subsequently, FlashMask wang2024flashmask utilizes the spatial continuity of non-zero mask values and proposes to only record the start and end indices of mask segments, achieving fast computation under rich mask types. Although the above works can speed up a wide range of mask types, as attention bias is usually continuous and dense, these sparsity-based methods cannot be applied to attention bias.
Recently, FlexAttention dong2024flex , taking advantage of deep learning compiler techniques, has been released in PyTorch 2.5, which supports a wide range of masks (e.g. causal mask achiam2023gpt ) and biases (e.g. ALiBi press2022train ) by pre-compiling element-wise operations. However, since many on-chip calculations are less optimized than matrix multiplications, FlexAttention cannot achieve a perfect speedup and fails to support data-dependent dynamic biases. In contrast, FlashBias can take full advantage of extremely optimized matrix multiplications and is also applicable to complex dynamic biases.
3 Method
As aforementioned, attention with bias is essential for model learning, while its natively dense and quadratic tensor shape results in a serious IO bottleneck. Inspired by the compressed sensing theory donoho2006compressed ; candes2010power , we verify that many well-established attention biases are inherently of low rank, paving the way for fast computation of attention with bias. Specifically, we dive into the computation of FlashAttention and theoretically prove that the optimal efficiency is inherently determined by the rank of attention weights and bias terms. Further, we present FlashBias based on the low-rank decomposition, which achieves optimal efficiency with natural adaptation for on-chip computing.
3.1 Rethinking FlashAttention Computation
We begin by analyzing the theoretical basis of FlashAttention’s speedup (without bias or mask) on dense and continuous attention weights. Our analysis yields that the rank of dense matrices, such as the dot-product attention weight $\mathbf{s}=\mathbf{q}\mathbf{k}^{\top}\in\mathbb{R}^{N\times M}$ and the bias matrix $\mathbf{b}\in\mathbb{R}^{N\times M}$, inherently decides the IO cost, which is formally stated as follows. All the proofs can be found in Appendix A.
Theorem 3.1 (FlashAttention computation benefits from low rank).
Suppose $N=M$ and let $R$ be the rank of dot-product attention weight $\mathbf{s}$, $C=\alpha N$ be the channel dimension with constant $\alpha$ and sequence length $N$, $S$ be the size of SRAM with $S=\beta NC$ and $\frac{1}{N}\leq\beta\leq 1$. Then, 1) the HBM access of FlashAttention is $\Theta\left((1+\frac{1}{\alpha})\beta\right)$ times smaller than the standard attention, and 2) $\alpha\geq\frac{R}{N}$.
As demonstrated in Theorem 3.1, the speedup of FlashAttention is proportional to $\beta$ (SRAM size) and inversely proportional to $\alpha$ (channel dimension). If we consider the attention weight $\mathbf{s}$ as a low-rank matrix, the optimal speedup of FlashAttention is obtained by reducing the channel dimension to $R$, i.e. $\alpha=\frac{R}{N}$. The same technique is also used in DeepSeek-v3 liu2024deepseek as Multi-Head Latent Attention, which reduces the channel dimension by projecting $\mathbf{q},\mathbf{k},\mathbf{v}$ into a small latent space for acceleration.
Theorem 3.2 (Compressed sensing complexity of low-rank dense matrix candes2010power ).
Given a $N\times N$ dense matrix with rank $R$, the theoretically optimal compressed tensor is of storage complexity $\Theta(NR)$.
Theorem 3.2 demonstrates that the optimal storage of bias is linearly related to its rank, highlighting the essentiality of the low-rank property. Further, integrating with prior analyses of HBM access in exact attention dao2023flashattention , we can derive the IO complexity property of HBM access in attention with bias.
Corollary 3.3 (A “lower bound” for HBM access of attention with bias).
Given $\mathbf{q}\in\mathbb{R}^{N\times C},\mathbf{k},\mathbf{v}\in\mathbb{R}^{M\times C}$, bias $\mathbf{b}$ of rank $R$ and SRAM of size $S$ where $(C+R)\leq S\leq N(C+R)$ , there does not exist an algorithm to compute exact attention with bias through $o\big{(}\frac{NM(C^{2}+R^{2})}{S}\big{)}$ HBM access for all $S\in[(C+R),N(C+R)]$ Here $o(\ast)$ represents the strict asymptotic upper bound.
Figure 2: Comparison between FlashAttention and FlashBias. FlashBias utilizes low-rank decomposition to bypass the read of the whole bias matrix, successfully avoiding the quadratic IO overload.
3.2 FlashBias
Inspired by the above theoretical results, we present FlashBias based on low-rank decomposition techniques, with novel design to utilize the low-rank property of attention bias to reduce HBM access.
Overall design
As shown in Figure 2, instead of block-wise reading bias, FlashBias replaces the quadratic bias matrix $\mathbf{b}\in\mathbb{R}^{N\times M}$ as two factor tensors. Specifically, considering a bias matrix calculated by $\mathbf{b}=f(\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}})$, where $\mathbf{x}_{\mathbf{q}}\in\mathbb{R}^{N\times C^{\prime}},\mathbf{x}_{\mathbf{k}}\in\mathbb{R}^{M\times C^{\prime}}$ represent the source information for generating the bias, which is set as spatial position of each pixel in Swin Transformer liu2021Swin and the representation of protein residues in AlphaFold abramson2024accurate , if there exist factor functions $\phi_{\mathbf{q}},\phi_{\mathbf{k}}$ satisfying:
$\begin{split}f(\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}})=\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q}})\phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k}})^{\top},\ \ \phi_{\mathbf{q}},\phi_{\mathbf{k}}:\mathbb{R}^{C^{\prime}}\to\mathbb{R}^{R}.\end{split}$ (2)
The computation of attention with bias can be equivalently formalized as follows:
$\begin{split}\mathbf{o}=\operatorname{softmax}(\frac{\mathbf{q}\mathbf{k}^{\top}}{\sqrt{C}}+\mathbf{b})\mathbf{v}=\operatorname{softmax}\big{(}\frac{\big{[}\mathbf{q}|\sqrt{C}\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q}})\big{]}\left[\mathbf{k}|\phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k}})\right]^{\top}}{\sqrt{C}}\big{)}\mathbf{v},\end{split}$ (3)
where [|][$\ast$|$\ast$][ ∗ | ∗ ] denotes the concatenation operation along the channel dimension. Notably, this design significantly reduces the storage cost for the attention bias from $\mathcal{O}(NM)$ to $\mathcal{O}\left((N+M)R\right)$. Although its design will require recalculating the bias weight, this computation is just a simple matrix multiplication of $\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q}})\phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k}})^{\top}$, an operation that has been extremely optimized on modern GPUs.
Such a simple design is broadly applicable to a wide range of variants for attention with bias. In practice, we implement it through three concrete instantiations for $\phi_{\mathbf{q}},\phi_{\mathbf{k}}$, as shown in Table 3.2.
Table 1: The computation of FlashBias for widely-used attention biases, which includes three different types: (a) Exact decomposition by finding exact $\phi_{\mathbf{q}},\phi_{\mathbf{k}}$, (b) SVD decomposition for cases using model parameter as bias, (c) Neural decomposition for using model representation as dynamic bias.
Domain Bias / Model Type
Language ALiBi press2022train (a)
Vision Swin Trans. liu2021Swin (b)
Spatial Distance (a)
Science
Pangu-Weather bi2023accurate
AlphaFold abramson2024accurate (c)
Exact decomposition
We find that some well-established attention biases can be directly decomposed into factor functions, enabling fast and exact computation. Here are representative cases.
Example 3.4 (ALiBi press2022train in language models).
Given $\mathbf{x}_{\mathbf{q}}=[1,\cdots,N],\mathbf{x}_{\mathbf{k}}=[1,\cdots,M]$, the ALiBi bias is calculated as $f(\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j})=i-j$, which can be directly decomposed into a low-rank formalization by defining $\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q},i})=[1,i]$ and $\phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k},j})=[-j,1]$, corresponding to the case $R=2$. The original ALiBi also involves a causal mask, while we only focus on the bias term here.
Example 3.5 (Spatial distance in scientific problems).
Transformers has been used as surrogate models for PDE solving wu2024Transolver , especially for aerodynamic simulation. It is critical to introduce spatial distance to guide attention learning among massive computational points. Let $\mathbf{x}_{\mathbf{q}}=\mathbf{x}_{\mathbf{k}}\in\mathbb{R}^{N\times 3}$ record the 3D spatial positions of $N$ computation points, where $\mathbf{x}_{\mathbf{q},i}\in\mathbb{R}^{3}$ is the position of $i$-th point. For the spatial distance $f(\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j})=\|\mathbf{x}_{\mathbf{q},i}-\mathbf{x}_{\mathbf{k},j}\|_{2}^{2}$, it can be exactly decomposed as:
$\begin{split}\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q},i})&=[\mathbf{x}_{\mathbf{q},i,0}^{2},1,-2\mathbf{x}_{\mathbf{q},i,0},\mathbf{x}_{\mathbf{q},i,1}^{2},1,-2\mathbf{x}_{\mathbf{q},i,1},\mathbf{x}_{\mathbf{q},i,2}^{2},1,-2\mathbf{x}_{\mathbf{q},i,2}],\\ \phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k},j})&=[1,\mathbf{x}_{\mathbf{k},j,0}^{2},\mathbf{x}_{\mathbf{k},j,0},1,\mathbf{x}_{\mathbf{k},j,1}^{2},\mathbf{x}_{\mathbf{k},j,1},1,\mathbf{x}_{\mathbf{k},j,2}^{2},\mathbf{x}_{\mathbf{k},j,2}].\end{split}$ (4)
SVD decomposition
Some models such as Swin Transformer liu2021Swin and Pangu-Weather bi2023accurate adopt the learnable model parameters for relative position encoding. Specifically, each bias term in their model is a $N\times M$ matrix of model parameters. As this type of bias is fixed once the model has been well trained, it is convenient to conduct Singular Value Decomposition (SVD) klema1980singular for low-rank decomposition of these parameters. In practice, we precompute SVD once offline, incurring negligible runtime overhead. The resulting decomposed factor tensors can then be utilized to accelerate the subsequent inference process, thanks to their low-rank nature.
Neural decomposition
The two mechanisms discussed above apply to static bias terms. Beyond these, some models introduce more complex or dynamically generated biases. For example, in AlphaFold abramson2024accurate , the bias term is projected from the intermediate pair representations. In this case, its low-rank decomposition cannot be explicitly or exactly derived. Also, due to the data-dependent property, SVD decomposition needs to be conducted for bias at every inference. To accelerate these complex biases, we present a neural approximation version of FlashBias, which employs two lightweight neural networks $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}:\mathbb{R}^{C^{\prime}}\to\mathbb{R}^{R}$ to approximate factor functions $\phi_{\mathbf{q}}$ and $\phi_{\mathbf{k}}$, which is supervised by the following objective function:
$\begin{split}\min_{\theta_{1},\theta_{2}}\mathcal{L}(\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}})=\|\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{q}})\widehat{\phi}_{\mathbf{k},\theta_{2}}(\mathbf{x}_{\mathbf{k}})^{\top}-f(\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}})\|_{2}^{2}.\end{split}$ (5)
Here $\theta_{1},\theta_{2}$ are learnable parameters, which can be optimized by fine-tuning $\theta_{1},\theta_{2}$ on the training set. Note that FlashBias attempts to completely replace the original bias term, a design fundamentally different from LoRA hu2022lora which learns an additive term to the pretrained model parameters. Similar to the SVD decomposition version, once these two lightweight neural networks $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ have been well optimized, they can be directly applied to all future inference.
Corollary 3.6 (HBM access of FlashBias).
Given $\mathbf{q}\in\mathbb{R}^{N\times C},\mathbf{k},\mathbf{v}\in\mathbb{R}^{M\times C}$, rank-$R$ bias $\mathbf{b}$ and size-$S$ SRAM ( $C\leq S\leq NC$ ), the HBM access complexity of FlashBias is $\Theta\big{(}\frac{NM(C^{2}+R^{2})}{S}\big{)}$.
Remark 3.7 (Trade-off between approximation accuracy and efficiency).
Although SVD and neural decomposition will introduce approximation error, owing to the native low-rank property of attention bias in modern Transformers, it will not affect the final performance by setting $R$ as a reasonable value. For example, in some layers in Swin Transformer, $R=32$ can maintain over 99% energy of the original bias with shape $576\times 576$. More case studies are included in Appendix C and D.
Example 3.8 (Comparison with FlashAttention).
For FlashAttention with bias, its HBM access is $\Theta(\frac{NMC^{2}}{S}+NM)$ dao2023flashattention . Given a regular setting of Transformers and GPU ($C=64$, $S$ is 100KB) and supposing $R=64$, $N,M\gg C,R$, then HBM IO of FlashBias is $\approx 6$, smaller than FlashAttention.
Speed up training
In FlashBias, the exact decomposition version can be directly applied to both training and inference phases. As for SVD or neural decomposition, their computations are based on a pretrained model; thereby, the inference speedup is obvious. Note that for large-scale pretrained models, such as AlphaFold abramson2024accurate , the inference speedup is already sufficiently valuable. Going further, it is also possible to speed up the training phase with these two types of biases. Specifically, for the SVD type, we can replace the $N\times M$ model parameter bias with $N\times R$ and $M\times R$ model parameter tensors at the initialization phase. As for the neural type, instead of using the whole bias matrix, it can also be replaced with two lightweight neural networks when defining the model architecture.
4 Experiments
Table 2: Summary of experiments and base models.
Base Model Bias Type Phase
Plain Transformer NIPS2017_3f5ee243 Static Train & Inference
GPT-2 radford2019language Static Train & Inference
Swin Trans. liu2022swin Learnable Inference
PDE Solver wu2024Transolver Learnable Train & Inference
AlphaFold 3 abramson2024accurate Learnable Inference
As summarized in Table 2, we extensively test FlashBias on Transformers for language modeling, computer vision and scientific problems, which can significantly speed up model efficiency at both training and test phases. This section will first present an overall efficiency comparison among different techniques and then integrate FlashBias to different domain-specific Transformers with various bias types. All the experiments were performed in PyTorch 2.5.0 Paszke2019PyTorchAI on a single A100 40GB GPU.
4.1 Overall Comparison
Setups
To give a clear comparison among FlashBias, vanilla FlashAttention with Bias dao2023flashattention and the latest FlexAttention dong2024flex , we make a comprehensive efficiency evaluation based on a plain Transformer NIPS2017_3f5ee243 , which consists of 8 layers. Each Transformer layer involves a feedforward network with 1024 intermediate hidden channels and attention with 512 hidden channels, 8 heads, as well as a static bias matrix of shape #$\#\operatorname{heads}\times N\times N$. All the metrics are recorded with a batch size of 1.
Figure 3: Efficiency comparison among FlashBias, FlashAttention w/ Bias and FlexAttention dong2024flex . Here “Pure FlashAttention” refers to canonical FlashAttention without a bias term, which can be viewed as an upper bound of computation efficiency. Dotted lines indicate out-of-memory situations.
Figure 4: Efficiency ratio over “Pure FlashAttention”, which is calculated by Method EfficiencyPure FlashAttention $\frac{\text{Method Efficiency}}{\text{Pure FlashAttention Efficiency}}$.
Results
As shown in Figure 3 and 4, we can find that FlashBias (red lines) consistently outperforms FlashAttention with Bias dao2023flashattention and FlexAttention dong2024flex , demonstrating the effectiveness of our design.
According to Figure 3(a-b), when sequence length $N=16384$, FlashBias can significantly reduce GPU memory usage, which is 5$\times$ smaller than FlexAttention and FlashAttention with Bias during training and 10$\times$ smaller during inference. This memory benefit can also be theoretically justified by Theorem 3.2. As for training and inference running time, FlashBias presents 18.6% and 44% speedup compared to vanilla usage of FlashAttention and is also better than FlexAttention in long sequences.
It is worth noticing that, although FlexAttention dong2024flex , which optimizes FlashAttention with deep learning compiler techniques, can significantly boost the running time compared to vanilla FlashAttention with Bias (Figure 4 (c-d) orange lines), it still cannot reduce GPU memory due to the quadratic storage cost of the bias matrix ((a-b) orange lines). In addition, due to the HBM access overload of the bias matrix, it falls short in running time when inputting long sequences. These observations further demonstrate that the low-rank property utilized in FlashBias is the key to efficiency.
4.2 Large Language Model
Setups
Here we employ FlashBias to speed up ALiBi press2022train , which is a well-known bias in language modeling and has been proven better than rotary position embedding su2024roformer in handling input length extrapolation. We follow the configuration in GPT-2 radford2019language and evaluate based on a large language model with ALiBi bias, which contains 48 decoder-only Transformer layers, 1.5B parameters in total. Each layer contains attention with 1600 channels and 50 heads, as well as a feedforward network with 6400 hidden dimensions. Beyond large model size, this task also involves the causal mask.
Implementation
As stated in Example 3.4, we can adopt the exact decomposition for ALiBi bias, where $R$ is equal to 2 and the result of FlashBias is exactly equivalent to the original computation. Since attention masking has been well-optimized in FlashAttention dao2023flashattention and FlexAttention dong2024flex , we directly integrate them with our method to support the decoder-only Transformer. This combination also demonstrates that FlashBias is orthogonal to attention-masking speed-up techniques.
[…]
Method Training Inference
Time(s) $\Delta$ Time(s) $\Delta$
Pure Causal FlashAttention 119.3 - 38.77 -
FlashAttention with Bias 124.3 5.0 40.32 1.55
FlashBias (Ours) 121.6 2.3 39.26 0.49
Pure Causal FlexAttention 119.0 - 38.76 -
FlexAttention with Bias 122.4 3.4 40.03 1.27
FlashBias (Ours) 121.5 2.5 39.78 1.02
Results
As mentioned above, since this task requires causal attention, we test FlashBias by integrating it with the attention mask speedup techniques in FlashAttention and FlexAttention. Results in Table 3 demonstrate that FlashBias still outperforms FlashAttention and FlexAttention in processing the ALiBi bias term. Specifically, in processing bias, FlashBias reduces over 50% (5.0$\to$2.3) time cost of FlashAttention and over 25% (3.4$\to$2.5) time cost of FlexAttention.
Note that ALiBi bias is relatively simple, enabling the compiler-based FlexAttention with a favorable speedup. However, it may degenerate in handling more complex bias matrices, such as the bias in Swin Transformer liu2022swin in the next section.
4.3 Vision Transformer
Setups
We test FlashBias on the image classification task based on Swin Transformer v2 liu2022swin . Specifically, we adopt the open-source SwinV2-B model [^1] Model and training configurations of SwinV2-B.. This model contains 24 layers with input resolution as $384\times 384$ and window size as $24\times 24$, thus, the sequence length of its WindowAttention is 576. WindowAttention in every layer contains a relative position bias with size #$\#\operatorname{heads}\times 576\times 576$, which is set as a learnable model parameter. We attempt to speed up the WindowAttention computation with FlashBias. All the efficiency metrics are evaluated under a batch size of 64.
Implementation
Since the biases are set as learnable parameters, we directly obtain them from a pretrained model and adopt the SVD decomposition to generate decomposed factor tensors. However, we observed that not all heads present low-rank bias matrices, as shown in Figure 5. Thus, we propose to split the heads into two subsets according to the biases’ ranks. For the subset with large ranks, we still adopt the vanilla FlashAttention with Bias, while adopting FlashBias for the low-rank subset.
Figure 5: Visualization of bias matrix in SwinV2-B. $R$ can maintain 99.5% energy of the bias matrix.
Table 4: Experiment of SwinV2-B on ImageNet-1K. #Time and #Mem correspond to inference efficiency on A100 per batch. Offline calculation of SVD for all biases takes 4.79s.
Method Acc@1 Acc@5 Time(s) Mem(MB)
Official Code 87.144% 98.232% 0.473 12829
Pure FlashAttention 9.376% 19.234% 0.180 4330
FlashAttention with Bias
87.142% 98.232% 0.230 10957
FlexAttention dong2024flex 87.142% 98.232% 2.885 25986
INT8 PTQ 86.46% Around 22% speed up
FlashBias (Ours) 87.126% 98.234% 0.190 5484
Results
In Table 4, we can find a serious performance drop in “Pure FlashAttention” (without bias, Acc@1 87$\to$9%), which verifies the essentiality of bias and also highlights the importance of speedup attention with bias. Further, compared to the official implementation, FlashBias can reduce 60% running time and 57% memory cost, which is valuable for real-time development. More importantly, FlashBias will only slightly affect the final performance, where the top-1 accuracy drops less than 0.02% and the fluctuation of the top-5 accuracy (0.002%) is within the standard deviation. In particular, even using well-established quantization techniques still will bring a 0.64% top-1 accuracy drop for 22% speedup according to the FasterTransformer document [^2] FasterTransformer released by NVIDIA.. This comparison demonstrates that the low-rank property is a principal basis for fast computation.
Also, FlexAttention seriously degenerates in SwinV2-B. This is because, unlike experiments in Section 4.1, Swin Transformer’s bias matrices are different in value and shape among different layers, which requires recompilation each time. This issue has also been mentioned by FlexAttention’s author as “If you are adding a [B, H, N, N] bias tensor, then you honestly shouldn’t be using FlexAttention [^3] Discussion about FlexAttention with dynamic bias matrix..
4.4 Scientific Deep Models
In addition to language and vision tasks, scientific problems usually involve rich domain-specific prior knowledge; thereby, attention bias also widely exists in scientific Transformers. Here we evaluate FlashBias in two representative models: Transformer-based PDE solvers wu2024Transolver and AlphaFold 3 abramson2024accurate .
Transformer-based PDE solvers
Attention mechanism has been proven equivalent to a Monte-Carlo approximation of the integral operator jmlr_operator , justifying its theoretical foundation for solving partial differential equations (PDEs). However, in processing complex geometries, the attention mechanism may fall in perceiving the 3D space, encouraging the utilization of spatial distance prior. Here we follow the driving car simulation task in wu2024Transolver , whose input is the position of computation mesh points and output is the physics quantities on these points. We test FlashBias on an 8-layer Transformer solver with a 3D distance bias described in Example 3.5. Each layer contains attention with 128 hidden dimensions and 8 heads, as well as a feedforward network with 256 hidden channels.
To approximate the adaptive mesh in numerical methods solin2005partial , we include a token-wise learnable weight $\alpha_{i}$ for the 3D distance bias in each head of every layer, i.e. $f(\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j})=\alpha_{i}\|\mathbf{x}_{\mathbf{q},i}-\mathbf{x}_{\mathbf{k},j}\|_{2}^{2}$. Unlike bias discussed in ALiBi press2022train or SwinV2 liu2021Swin , the learnable weights require the training process to record the gradient of the bias matrix, posing a challenging efficiency issue in backpropagation.
Table 5: Experiments of Transformer PDE solvers. The efficiency metrics are recorded under a batch size of 1 in the format of #Mem (GB) / #Time (s/100iters). Accuracy comparisons are in Appendix E.
Method (Learnable Bias) Training Phase Inference Phase
$8192$ $16384$ $32186$ $8192$ $16384$ $32186$
FlashAttention 12.8 / 15.4 OOM OOM 4.54 / 5.46 15.3 / 21.2 OOM
FlexAttention Not supported in current version 21.9 / 184.0 OOM OOM
FlashBias (Ours) 1.46 / 4.54 2.02 / 14.7 2.97 / 51.1 0.98 / 1.22 1.03 / 3.48 1.13 / 12.7
Results
FlashBias is the only method that can support training of the Transformer solver on 32186 points (Table 5), which also presents a significant memory and running time advantage compared with other methods. Notably, FlashAttention and FlexAttention cannot support learnable bias training well, as they need to record dense gradient matrices, further highlighting the practicality of FlashBias.
AlphaFold 3 for protein folding
As a significant progress of AI for science, AlphaFold 3 abramson2024accurate employs an elaborate architecture. Specifically, its core design, Pairformer, contains 144 attention blocks, all of which contain the bias term of pair representations. Thus, its speedup is highly related to the fast computation of attention with bias. Since Alphafold 3 is not officially open-sourced, we follow a public high-quality reproduction, Proteinix chen2025protenix , which is implemented in PyTorch Paszke2019PyTorchAI .
After a detailed analysis of AlphaFold 3, we find that its efficiency bottleneck is triangle self-attention, which involves the bias matrix projected from intermediate pair representations, making it vary among different samples, layers and heads. To approximate this complex bias, we employ the neural decomposition version of FlashBias, whose inputs $\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}}$ are set as the combination of pair and single protein representation. We fine-tune neural basis functions for 2000 iterations on the PDB dataset. Since we only need to optimize the newly added parameters in $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ , this process only takes about 5 hours on a single A100 40GB GPU, then you can infer a new protein with FlashBias. For comparison, the full training process of AlphaFold 3 will take about 7 days on 128 A100 GPUs.
Table 6: Experiment of AlphaFold 3. The left part illustrates folding examples. Note that AlphaFold 3 is based on a diffusion model; thereby, slight differences are normal. The right table’s efficiency is tested on the one-time inference of Pairformer with 1218 protein residue tokens (PDB ID: 7wux).
Method Test Set PDB ID 7wux
pLLDDT Loss $\downarrow$
pTM $\uparrow$
Time(s)
Mem(GB)
Open-sourced Code 3.3724 0.9500 31.48 21.58
FlashAttention w/o Bias
4.3669 0.1713 18.81 15.76
FlashAttention w/ Bias
3.3724 0.9500 26.17 18.69
FlashBias (Ours) 3.3758 0.9498 20.89 10.37
Results
Table 4.4 shows that, compared to the public code, FlashBias can reduce the running time and GPU memory usage by 30% and 50%, respectively. This speedup is very close to the efficiency upper bound: FlashAttention w/o Bias. Although FlashBias is based on the neural decomposition, it will not affect the final performance, whose metric fluctuation is within the standard deviation. Beyond inference, FlashBias can also be a promising tool for training speedup (see Appendix C).
5 Conclusions
This paper focuses on the fast computation of attention with bias, which is an essential extension of the attention mechanism and is widely used in language, vision and scientific domains. After an in-depth analysis of FlashAttention, we notice that the optimal efficiency depends on the low-rank property of the attention weight. This further inspires us to present FlashBias based on the low-rank compressed sensing theory, where we also present three practical methods for low-rank decomposition of attention bias, achieving theoretically favorable efficiency. Experimentally, FlashBias can seamlessly support the fast computation of a wide range of famous Transformers without loss of accuracy.
Appendix A Proofs for Theorems in the Main Text
Due to the page limitation of the main text, we present the proofs for theorems in Section 3 here.
Proof of Theorem 3.1
This theorem can be proven based on the theoretical analyses in FlashAttention wang2024flashmask and the basic low-rank compressed sensing knowledge.
Proof.
According to FlashAttention wang2024flashmask , the theoretical complexity of HBM accesses in the standard attention and FlashAttention are:
$\begin{split}\operatorname{FlashAttention:}\ \text{IO}_{\text{flash}}=\Theta\left(\frac{N^{2}C^{2}}{S}\right),\ \operatorname{StandardAttention:}\ \text{IO}_{\text{standard}}=\Theta\left(NC+N^{2}\right),\end{split}$ (6)
where $\Theta$ represents the asymptotic tight bound. Given $C=\alpha N$ and $S=\beta NC$, the ratio of HBM access in standard attention over FlashAttention is:
$\frac{\text{IO}_{\text{flash}}}{\text{IO}_{\text{standard}}}=\Theta\left(\frac{(NC+N^{2})S}{{N^{2}C^{2}}}\right)=\Theta\left(\frac{(\alpha+1)\alpha\beta}{\alpha^{2}}\right)=\Theta\left(\beta(1+\frac{1}{\alpha})\right).$ (7)
Also, since $R$ is the rank of the attention weight $\mathbf{s}$, which is the dot-product of queries and keys, thus $R\leq C$. Further, we can derive that $\alpha\geq\frac{R}{N}$. ∎
Proof of Theorem 3.2
As cited in the main text, the least storage cost of an $N\times N$ dense matrix with rank $R$ is equal to $(2NR-R^{2})$, whose proof can be found in candes2010power . Since $R\leq N$, then
$NR\leq 2NR-R^{2}\leq 2NR.$ (8)
Thus, the optimal storage overload for an $N\times N$ matrix is equal to $\Theta(NR)$.
Proof of Corollary 3.3
This corollary is derived from Proposition 3. in FlashAttention dao2022flashattention . Specifically, consider the extreme case that $S=\Theta\left(N(C+R)\right)$, in this case, we have
$o\left(\frac{NM(C^{2}+R^{2})}{S}\right)=o\left(\frac{NM(C^{2}+R^{2})}{N(C+R)}\right)=o\left(M(C+R)\right).$ (9)
Since we have to read the $M\times C$ keys and $M\times R$ decomposed bias factor tensor from HBM, there does not exist an algorithm to finish the computation of attention with bias in $o\left(M(C+R)\right)$ access.
Proof of Corollary 3.6
This corollary can be derived from Theorem 2 in FlashAttention dao2022flashattention .
Proof.
As formalized in Eq. 3, FlashAttention needs to read queries, keys, values, as well as decomposed factor tensors of biases into the on-chip SRAM for computation.
FlashBias’s computation is based on the same tiling method as FlashAttention v2 dao2023flashattention . The $N\neq M$ case corresponds to the cross-attention, which is used for fusing additional information jaegle2021perceiver , thus we suppose $M\geq N$ here. Suppose that the computation splits queries into $T$ blocks. The overall HBM access complexity is $\Theta\left(N(C+R)+M(C+R)T\right)=\Theta\left(M(C+R)T\right)$. Let the block size of keys and values as $B_{\mathbf{k,v}}\times(C+R)$ and the block size of queries as $B_{\mathbf{q}}\times(C+R)$. Considering the size limitation of SRAM is $S$, following FlashAttention v2 dao2023flashattention , these hyperparameters are set as:
$B_{\mathbf{q}}=\Theta\left(\frac{S}{C+R}\right),\ \ B_{\mathbf{k,v}}=\Theta\left(\min\left(\frac{S}{C+R},\frac{S}{B_{\mathbf{q}}}\right)\right)=\Theta\left(\min\left(\frac{S}{C+R},C+R\right)\right).$ (10)
Then, the number of blocks $T=\frac{N}{B_{\mathbf{q}}}=\Theta(\frac{N(C+R)}{S})$. Thus, the overall HBM access complexity is
$\Theta\left(\frac{NM(C+R)^{2}}{S}\right)=\Theta\left(\frac{NM(C^{2}+R^{2})}{S}\right).$ (11)
Recall Corollary 3.3, we can find that the above theoretical complexity is quite well-optimized as there does not exist an algorithm with $o\left(\frac{NM(C^{2}+R^{2})}{S}\right)$ complexity. ∎
Calculation in Example 3.8
Considering $N,M\gg C,R$, it is easy to calculate the HBM access ratio between FlashAttention and FlashBias as follows:
$\frac{(1+\frac{C^{2}}{S})S}{C^{2}+R^{2}}\approx\frac{(1+\frac{2\times 64^{2}}{100\times 1024})100\times 1024}{2(64^{2}+64^{2})}\approx 6,$ (12)
where we consider the half-precision float, whose storage cost is equal to 2B.
Appendix B Speedup of Pangu-Weather bi2023accurate
Pangu-Weather bi2023accurate is a significant step in adopting Transformers for global weather forecasting. Specifically, its backbone is a 3D Swin Transformer with a hierarchical structure, which contains two different scales. Its speedup is similar to our experiments in Section 4.3. As listed in Table 3.2, we consider to speedup this model based on the SVD decomposition version of FlashBias.
Setups
Since Pangu-Weather is based on the 3D window (with shape $2\times 6\times 12$), its bias for relative position encoding is slightly different from Swin Transformer. Especially, its bias is in the shape of #$\#\operatorname{num}\times\#\operatorname{heads}\times 144\times 144$ for each block, where #$\#\operatorname{num}$ represents the number of 3D windows. According to meteorological knowledge, different longitudes share the same bias.
Implementations
Since Pangu-Weather does not provide accessible model weights or code, our experiments are based on an open-sourced PyTorch reproduction [^4] https://github.com/zhaoshan2/pangu-pytorch. All the other implementations are the same as the descriptions in Section 4.3. Notably, we find that only relative position biases in the fine scale are low-rank. Thus, we only apply FlashBias to the four 3D Swin layers in the fine scale, where we set $R=56$ to maintain 99% energy of the original bias matrix. As discussed in Section 4.3, FlexAttention fails in processing such dynamic bias; we didn’t compare with it in this task. Since the whole ERA5 data is over 150TB, we only test the model based on 100 samples in 2024.
Table 7: Experiments of Pangu-Weather bi2023accurate on ERA5 hersbach2020era5 . Output difference measures of the L2 distance between the outputs of FlashBias and the official code, averaged from 100 different inputs. This difference is calculated on the z-score normalized patro2015normalization model outputs to balance diverse meteorological factors.
Method
Output Difference
Time(
s/100iters
Mem(MB)
Open-sourced Code - 98.022 26552
FlashAttention w/o bias 0.0128 74.089 12141
FlashAttention w/ bias - 79.649 13186
FlashBias (Ours) 0.0003 76.779 12222
Results
As presented in Table 7, FlashBias can also speedup Pangu-Weather. Compared to the open-sourced code, FlashBias reduces over 20% running time and over 50% GPU memory usage. However, due to the limited sequence length ($N=144$ in this case), the running time speedup is not as considerable as SwinV2. Just as plotted in Figure 3 and 4, FlashBias will bring more significant speedup in running time and GPU memory for long sequences. As improving the spatial resolution of forecasting is a golden problem in weather prediction wu2023interpretable , we believe that FlashBias has a strong advantage in supporting future research on higher-resolution reanalysis data.
Appendix C More Results in AlphaFold 3 abramson2024accurate
Potential in speeding up training
In the main text, we only evaluate FlashBias during the inference phase. Going further, as mentioned at the end of Section 3.2, if we directly replace the bias matrix in AlphaFold 3 abramson2024accurate with two decomposed factor functions, we can also accelerate the training process. As shown in Table 8, FlashBias can save around 9% running time and 30% GPU memory usage.
Table 8: Experiments of FlashBias in accelerating the training process of AlphaFold 3, where the default setting is to crop all the residue sequence into 384 tokens.
Method Running Time (s / 10 iters) GPU Memory (GB)
Open-sourced Code 153.272 23.552
FlashAttention with bias 150.058 22.740
FlashBias (Ours) 140.281 16.260
However, since the complete training of AlphaFold 3 will require around 128 GPUs for 7 days, considering the resource limitation, we would like to leave the verification of the final performance of FlashBias-accelerated AlphaFold 3 as our future work.
Analysis of pair representation bias
To give a clear illustration of the learned neural decomposition, we also plot the original pair representation bias and FlashBias approximated bias in Figure 6. From the visualization, we find that neural decomposition can give a relatively accurate estimation of the bias matrix, which performs well in capturing the “texture” of the original bias matrix. In addition, it is also observed that neural decomposition is not completely perfect in reconstructing the diagonal weights. Despite this deficiency, FlashBias still maintains the original accuracy of AlphaFold3, as presented in Table 4.4 of the main text. This may benefit from the dot-product self-attention mechanism and residual connection, which can give a robust and dominating weight for self-relation modeling.
Figure 6: Comparison between neural decomposed factor tensors’ multiplication and original bias. Here we visualize the bias of 7r6r (245 residues) and 7pzb (600 residues) in the first layer of Pairformer, which contains four heads. Their concrete 3D folding structure is plotted in Table 4.4. We also mark the rank value that can maintain 99% energy of the original bias in the above figure.
Appendix D More Results in Swin Transformer V2 liu2022swin
Figure 7: Maximum rank value of the split two subsets across 24 layers in SwinV2-B.
Statistics of Swin Transformer V2 Bias
As described in Section 4.3, we split the bias matrix of every layer into two subsets according to the rank value that can maintain 95% energy. Here, each layer contains a bias matrix with shape #$\#\operatorname{heads}\times 576\times 576$ and each subset is of shape #$\frac{\#\operatorname{heads}}{2}\times 576\times 576$. As shown in Figure 7, the maximum rank value of the low-rank subset is smaller than 100 in 23 layers (except the third layer). Based on these statistical data, we set $R$ in FlashBias as [32, 32, 280, 40, 48, 40, 40, 88, 80, 32, 64, 32, 32, 32, 32, 32, 32, 88, 32, 32, 32, 32, 32, 32] for the low-rank subset in each layer. Especially, we set $R\equiv 0\pmod{8}$ following FlashAttention v2 dao2023flashattention to facilitate block-wise reading.
SVD decomposition visualization
In Figure 5 of the main text, we visualize the original bias matrix at the 15th layer with 16 heads. To deliver an intuitive understanding, we also take this layer’s bias as an example and compare it with the SVD decomposed factor tensors’ multiplication. Since we only adopt SVD decomposition on half heads for speedup, we only plot 8 heads here. As shown in Figure 8, SVD decomposed factor tensors can accurately reconstruct the original bias matrix.
Figure 8: Comparison between SVD decomposed factor tensors’ multiplication and original bias.
Appendix E More Results in Transformer PDE Solver
In Table 5, we only present the efficiency comparison. Further, to demonstrate the performance benefits brought by spatial distance bias, we also include the performance metric in Table 9. With spatial distance bias, the error of the estimated drag coefficient can be reduced by 65.3%, which is a significant progress in industrial design. This further confirms the importance of attention with bias.
Table 9: Performance comparison among attention w/o bias and w/ bias in PDE solving. The relative L2 of surface pressure and surrounding velocity is recorded. We also calculate the drag coefficient $C_{D}$ based on model-predicted physics fields, whose relative L2 w.r.t. ground truth is also included.
Method (Sequence Length $N=32186$) Surface Pressure Error Surrounding Velocity Error $C_{D}$ Error
Pure attention without spatial distance bias 0.0838 0.0278 0.0173
FlashAttention with spatial distance bias OOM OOM OOM
FlashBias with spatial distance bias 0.0706 0.0201 0.0113
Relative Promotion 15.7% 27.7% 65.3%
Appendix F Generalization for Diverse Biases
In FlashBias, we present a neural decomposition version to fit complex and dynamic biases, which has been tested in speeding up AlphaFold 3 in Section 4.4. To further demonstrate the expressive capability of neural decomposition, in this section, we will train neural factor functions $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ to approximate more diverse biases, which can be meaningful for scientific tasks.
Figure 9: Adopting neural decomposition techniques in FlashBias for more diverse biases.
Gravity bias
Many phenomena are inherently governed by underlying physical forces, where gravity is one of the basic factors. Accurately approximating the gravity force is essential for the modeling of planetary motion russell1964kepler or electronic simulation elliott1993electromagnetics . Thus, we consider introducing the gravity bias into the attention mechanism. Specifically, this bias can be formalized as follows:
$f(\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j})=\frac{1}{\|\mathbf{x}_{\mathbf{q},i}-\mathbf{x}_{\mathbf{k},j}\|_{2}^{2}},$ (13)
where $\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j}$ denotes the spatial positions of the $i$-th and $j$-th objects. We train $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ based on randomly sampled points from $[0,1]\times[0,1]$ in the 2D space. Since the bias is inversely proportional to spatial distance, we further add 0.01 to the diagonal bias for numerical stability.
Spherical distance
When analyzing atmospheric circulation, it is intuitive to consider it as the dynamics on a spherical surface. Therefore, previous research has attempted to introduce spherical Fourier analysis into the model design bonev2023spherical . Another alternative approach is to add a spherical distance bias to attention in Transformer-based models. Thus, we also consider the spherical distance bias:
$f(\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j})=2\cdot\arcsin\left(\sqrt{\sin^{2}(\frac{\mathbf{x}_{\mathbf{q},i,0}-\mathbf{x}_{\mathbf{k},j,0}}{2})+\cos\mathbf{x}_{\mathbf{q},i,0}\cos\mathbf{x}_{\mathbf{k},j,0}\sin^{2}(\frac{\mathbf{x}_{\mathbf{q},i,1}-\mathbf{x}_{\mathbf{k},j,1}}{2})}\right),$ (14)
where $\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j}\in\mathbb{R}^{2}$ records the latitude and longitude of the $i$-th and $j$-th position respectively. Similarly, we also train $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ based on randomly sampled points in $[-\pi,\pi]\times[0,2\pi]$.
For the biases mentioned above, we set $R=32$ and $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ as three linear layers with in-between tanh activation functions. Then we optimize these model parameters with the Adam DBLP:journals/corr/KingmaB14 optimizer for 10000 iterations, which will take less than 30 seconds on an A100 GPU. As presented in Figure 9, neural decomposition performs very well in these two biases, especially for the spherical bias. As for gravity bias, since the numerical instability of the inverse proportion function, it is more difficult for optimization, while our method still captures the locality of the bias matrix.
The above experiments further testify the capability of FlashBias in accelerating broader scenarios.
Appendix G Implementation Details
As a supplement to the main text, we include more details of implementation here. All the experiments are conducted based on PyTorch 2.5.0 Paszke2019PyTorchAI and on a single A100 GPU with 144 CPU cores. All efficiency metrics are averaged from 1000 iterations. For example, some metrics are recorded as s/100iters or s/10iters, where we divide the running time of 1000 iterations by 10 or 100, respectively. In our experiments, all the algorithms’ efficiency performance is quite stable.
ALiBi in Large Language Model
Since the experiments for ALiBi speedup are based on directly replacing the ALiBi bias with the exact decomposition, FlashBias’s output results are completely equal to the original version. The training efficiency is tested under the Adam DBLP:journals/corr/KingmaB14 optimizer.
Relative position encoding in Swin Transformer
This part of the experiments strictly follows the official code in Swin Transformer V2 liu2022swin . As illustrated in Figure 7, we only apply FlashBias to the computation of half heads in each layer (except the third layer). To further reduce the memory addressing cost, we also permute the linear layer parameters to move the speedup heads to the first half dimension, including the linear layer for generating queries, keys, values and the linear layer for final output. This operation can also be finished offline and directly applied to the inference phase.
Spatial distance bias in PDE solver
In this experiment, we follow the code base provided here  [^5] https://github.com/thuml/Neural-Solver-Library, which includes the application of Transformers for the car design task. Specifically, the input contains the position of the pre-defined computation mesh and the output is the pressure and air velocity on these computation mesh points. In this task, we also adopt the exact decomposition for computation. The training phase is also based on the Adam DBLP:journals/corr/KingmaB14 optimizer with a batch size of 1.
Pair representation bias in AlphaFold 3
In AlphaFold 3 abramson2024accurate , we adopt the neural decomposition version. This version involves the training of newly added neural layers, whose finetuning configuration is listed in Table 10. Here we find that optimizing the factor functions for 2000 iterations can obtain a nearly converged version. This may be because the factor functions are token-wise, which means at every iteration, the model will receive $N$ (sequence length) samples for training. To sum up, these layers have been optimized with around 768000 samples after 2000 iterations. Also, to reduce the cumulative error, we only apply FlashBias to the first 16 Pairformer blocks.
Table 10: Configuration for finetuning factor functions $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$.
Part Configuration
Input $\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}}$ Sum of row and column in pair representation, with shape of $N\times 128$
Single representation at model beginning with shape of $N\times 449$
Output $\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{q}})$, $\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{k}})$ Decomposed factor tensors with shape of $N\times\#\operatorname{heads}\times 64$ (
#$\#\operatorname{heads}=4$
Model $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ Input Dim: 577; Hidden Dim: 256; Output Dim: 256 ($=4\times 64$)
Three linear layers with in-between tanh activation function
Initial learning rate: 0.001; Optimizer: Adam;
Training Learning rate decay: every 50 iterations, reduce to the origin’s 0.95
Overall steps: 2000 iterations
Dataset Train set: “weightedPDB_before2109_wopb_nometalc_0925”
Test set: “recentPDB_1536_sample384_0925”